a merchant has coffee worth $50 a pound that she wants to mix with 90 pounds of coffee worth $80 a pound to get a mixture that can be sold for $70 a pound. How many pounds of the $50 coffee should be used?
Answer by stanbon(75887)

a merchant has coffee worth $50 a pound that she wants to mix with 90 pounds of coffee worth $80 a pound to get a mixture that can be sold for $70 a pound. How many pounds of the $50 coffee should be used?Equation:
value + value = value
50x + 80*90 = 70(x+90)
-----
50x + 80*90 = 70x + 70*90
----
20x = 10*90
x = 45 lbs (amt. of $50 coffee needed)
===============================
Cheers,
stan H.